### v1

# The bracket of the exceptional Lie algebra E8

Abstract.
An explicit formula for the bracket of the exceptional simple Lie algebra E8 based on triality and oct-octonions is obtained, following the Barton-Sudbery description of E8.
Key words and phrases:
exceptional Lie algebra, octonions
2020 Mathematics Subject Classification:
17B25, 17A35, 22E60, 53C35
1. Introduction
The exceptional Lie groups and Lie algebras $\mathrm{G}_{2}$, $\mathrm{F}_{4}$, $\mathrm{E}_{6}$, $\mathrm{E}_{7}$, $\mathrm{E}_{8}$ were first described by Wilhelm Killing [10, 4]. The exceptional Lie groups have since made many appearances in geometry, topology and physics. However, to this day they still remain somewhat elusive and mysterious.
Of the five exceptional Lie algebras $\operatorname{\mathfrak{g}}_{2}$, $\operatorname{\mathfrak{f}}_{4}$, $\operatorname{\mathfrak{e}}_{6}$, $\operatorname{\mathfrak{e}}_{7}$, $\operatorname{\mathfrak{e}}_{8}$ the $248$-dimensional algebra $\operatorname{\mathfrak{e}}_{8}$ is the one with highest dimension, and in many regards also the most difficult to deal with. For instance, since the lowest dimensional non-trivial representation is $248$-dimensional, it is not an option to use a matrix representation of $\operatorname{\mathfrak{e}}_{8}$, an approach that is viable for $\operatorname{\mathfrak{g}}_{2}$, which has a $7$-dimensional faithful representation. For this and other reasons, the title of the article [8] refers to $\mathrm{E}_{8}$ as the “most exceptional Lie group” and it has been argued in [8], that $\mathrm{E}_{8}$ could be called “the Monster of Lie theory”. The difficulties that arise when computations with these objects are required, have led to computer-based approaches as in [13], where structure constants of $\operatorname{\mathfrak{e}}_{8}$ are determined.
While several descriptions and constructions of $\operatorname{\mathfrak{e}}_{8}$ can be found in the literature, see [1, 5, 14], so far, no explicit formula for its bracket appears to be known. In this paper, we give an explicit formula for the Lie bracket of the compact real form of $\operatorname{\mathfrak{e}}_{8}$, which drastically simplifies certain calculations with $\operatorname{\mathfrak{e}}_{8}$. Besides standard linear algebra operations, our formula (4.1) involves multiplication of Cayley numbers and the triality automorphism of $\operatorname{\operatorname{\mathfrak{so}}}(8)$. Our description is based on the decomposition
$\operatorname{\mathfrak{e}}_{8}\cong\operatorname{\operatorname{\mathfrak{so}}}(8)\oplus\operatorname{\operatorname{\mathfrak{so}}}(8)\oplus(\mathbb{O}\otimes\mathbb{O})^{3},$
which is called the Barton-Sudbery description of $\operatorname{\mathfrak{e}}_{8}$, cf. [3], by Baez [2]. In his article [2], Baez writes: “To emphasize the importance of triality, we can rewrite the Barton-Sudbery description of $\operatorname{\mathfrak{e}}_{8}$ as
$\operatorname{\mathfrak{e}}_{8}\cong\operatorname{\operatorname{\mathfrak{so}}}(8)\oplus\operatorname{\operatorname{\mathfrak{so}}}(8)\oplus(V_{8}\otimes V_{8})\oplus(S^{+}_{8}\otimes S^{+}_{8})\oplus(S^{-}_{8}\otimes S^{-}_{8})$
Here the Lie bracket is built from natural maps relating $\operatorname{\operatorname{\mathfrak{so}}}(8)$ and its three 8-dimensional irreducible representations. In particular, $\operatorname{\operatorname{\mathfrak{so}}}(8)\oplus\operatorname{\operatorname{\mathfrak{so}}}(8)$ is a Lie subalgebra, and the first copy of $\operatorname{\operatorname{\mathfrak{so}}}(8)$ acts on the first factor in $V_{8}\otimes V_{8}$, $S_{8}^{+}\otimes S_{8}^{+}$, and $S_{8}^{-}\otimes S_{8}^{-}$, while the second copy acts on the second factor in each of these. This has a pleasant resemblance to the triality description of $\operatorname{\mathfrak{f}}_{4}$ […]:
$\operatorname{\mathfrak{f}}_{4}\cong\operatorname{\operatorname{\mathfrak{so}}}(8)\oplus V_{8}\oplus S^{+}_{8}\oplus S^{-}_{8}.\hbox{''}$ ”
In the article [11], the author obtained an explicit bracket formula (2.5) for $\operatorname{\mathfrak{f}}_{4}$ in order to study a geometric problem on the Cayley hyperbolic plane. This formula is based on the above-mentioned triality description of $\operatorname{\mathfrak{f}}_{4}$. It turns out that there is a natural generalization (4.1) of this formula for the bracket of $\operatorname{\mathfrak{e}}_{8}$. We obtain it by replacing $\operatorname{\operatorname{\mathfrak{so}}}(8)$ with $\operatorname{\operatorname{\mathfrak{so}}}(8)\oplus\operatorname{\operatorname{\mathfrak{so}}}(8)$ and replacing the octonions $\mathbb{O}$ with the real algebra $\mathbb{O}\otimes\mathbb{O}$, the oct-octonions. The purpose of this paper is to prove that the formula (4.1) actually defines a Lie algebra, i.e., satisfies the Jacobi identity and this Lie algebra can be endowed with an invariant scalar product, showing that the resulting simple real Lie algebra is isomorphic to the compact real form of $\operatorname{\mathfrak{e}}_{8}$. This can also be regarded as a new existence proof for $\operatorname{\mathfrak{e}}_{8}$.
2. Preliminaries
Recall that the octonions, or Cayley numbers, are an $8$-dimensional non-associative and non-commutative real division algebra $\mathbb{O}$. We identify $\mathbb{O}$ with $\mathbb{R}^{8}$ by choosing a basis $e_{0},\dots,e_{7}$ of $\mathbb{O}$ such that $e:=e_{0}$ is the multiplicative identity, we have $e^{2}_{1}=\dots=e^{2}_{7}=-e$ and the following multiplication rules:
$\begin{array}[]{llll}e_{1}e_{2}=e_{3},&e_{1}e_{4}=e_{5},&e_{2}e_{4}=e_{6},&e_{3}e_{4}=e_{7},\\ e_{5}e_{3}=e_{6},&e_{6}e_{1}=e_{7},&e_{7}e_{2}=e_{5}.&\end{array}$ (2.1)
From these rules, the whole multiplication table of $\mathbb{O}$ can be obtained using the fact that the alternative laws $(xx)y=x(xy)$ and $(xy)y=x(yy)$ hold for all $x,y\in\mathbb{O}$. We define octonionic conjugation as the $\mathbb{R}$-linear map $\gamma\colon\mathbb{O}\to\mathbb{O},a\mapsto\bar{a}$, given by
$\gamma(e_{s})=\left\{\begin{array}[]{ll}e_{0},&\hbox{if~{}$s=0$;}\\ -e_{s},&\hbox{if~{}$s\geq 1$.}\end{array}\right.$
We identify the real numbers $\mathbb{R}$ with the subalgebra of $\mathbb{O}$ spanned by $e$ and define the real part of an octonion $x$ as the real number $\mathrm{Re}(x):=\frac{1}{2}(x+\bar{x})$ and the pure part as the octonion $\operatorname{Pu}(x):=\frac{1}{2}(x-\bar{x})$. The norm of an octonion $a$ is defined as $|a|=\sqrt{a\bar{a}}$. Recall that $\mathbb{O}$ is a normed algebra, i.e. we have $|ab|=|a|\,|b|$ for all $a,b\in\mathbb{O}$. For $u,v\in\mathbb{R}^{8}=\mathbb{O}$, we denote by $\langle u,v\rangle=\mathrm{Re}(u\bar{v})=u^{t}v$ the standard scalar product of $\mathbb{R}^{8}$.
Let $\operatorname{\operatorname{\mathfrak{so}}}(8)$ be the special orthogonal Lie algebra in $8$ dimensions, which consists of the real skew-symmetric $8\times 8$ matrices, where the bracket is given by the commutator of matrices. There is a map $\mathbb{O}\times\mathbb{O}\to\operatorname{\operatorname{\mathfrak{so}}}(8)$, $(x,y)\mapsto x\wedge y$, defined by
$x\wedge y:=xy^{t}-yx^{t},$ (2.2)
where, on the right hand side, the usual matrix multiplication is understood and $y^{t}$ is the row vector which is the transpose of $y$. The elements $e_{i}\wedge e_{j}$, $0\leq i<j\leq 7$, form a basis of $\operatorname{\operatorname{\mathfrak{so}}}(8)$.
For $a\in\mathbb{O}$, we define the $\mathbb{R}$-linear maps $L_{a},R_{a}\colon\mathbb{O}\to\mathbb{O}$ as $L_{a}(x)=ax$ and $R_{a}(x)=xa$, i.e. the left and right multiplication by $a$. Recall from [7], [12], or [11] that an automorphism of $\operatorname{\operatorname{\mathfrak{so}}}(8)$ of order three is given by the map
$\lambda(a\wedge b)=\tfrac{1}{2}L_{\bar{b}}\circ L_{a}\hbox{~{}for~{}$a\in\operatorname{Pu}(\mathbb{O}),b\in\mathbb{O}$}.$ (2.3)
Recall furthermore that $\lambda^{2}=\lambda\circ\lambda$ is given by
$\lambda^{2}(a\wedge b)=\tfrac{1}{2}R_{\bar{b}}\circ R_{a}\hbox{~{}for~{}$a\in\operatorname{Pu}(\mathbb{O}),b\in\mathbb{O}$}.$ (2.4)
We define another automorphism $\kappa\in\operatorname{Aut}(\operatorname{\operatorname{\mathfrak{so}}}(8))$ by
$\kappa(a\wedge b)=\bar{a}\wedge\bar{b},$
or, equivalently, $\kappa(A)x=\overline{A\bar{x}}$ for $x\in\mathbb{R}^{8}$. Then we have
$\kappa^{2}=\lambda^{3}=1\qquad\hbox{and}\qquad\kappa\circ\lambda^{2}=\lambda\circ\kappa,$
see [7], [12, §2, Thm. 2]. In [11] the following explicit expression was obtained for the Lie bracket of the simple compact Lie algebra $\operatorname{\mathfrak{f}}_{4}$.
Theorem 2.1.
The binary operation on $\operatorname{\operatorname{\mathfrak{so}}}(8)\times\mathbb{O}^{3}$ defined by
$[(A,u,v,w),(B,x,y,z)]=(C,r,s,t)$ (2.5)
where
$C$ $=AB-BA-4u\wedge x-4\lambda^{2}(v\wedge y)-4\lambda(w\wedge z),$
$r$ $=Ax-Bu+\overline{vz}-\overline{yw},$
$s$ $=\lambda(A)y-\lambda(B)v+\overline{wx}-\overline{zu},$
$t$ $=\lambda^{2}(A)z-\lambda^{2}(B)w+\overline{uy}-\overline{xv},$
is $\mathbb{R}$-bilinear, skew symmetric and satisfies the Jacobi identity. The real Lie algebra which is defined in this way is isomorphic to the simple compact Lie algebra $\operatorname{\mathfrak{f}}_{4}$.
[…]
3. Oct-octonions and $\operatorname{\operatorname{\mathfrak{so}}}(8)\oplus\operatorname{\operatorname{\mathfrak{so}}}(8)$
We define the oct-octonions to be the real algebra $\mathbb{O}\otimes\mathbb{O}$, where the multiplication is given by
$(a\otimes b)\ast(c\otimes d)=ac\otimes bd.$
Using the identification $\mathbb{O}=\mathbb{R}^{8}$, we may identify $\mathbb{O}\otimes\mathbb{O}$ with $\mathbb{R}(8)$, the space of real $8\times 8$ matrices. Indeed, viewing octonions as column vectors in $\mathbb{R}^{8}$, we make the following identification for $x,y\in\mathbb{O}$:
$x\otimes y=xy^{t},$ (3.1)
where, on the right hand side, the usual matrix multiplication is understood and $y^{t}$ is the row vector which is the transpose of $y$. We have
$(x\otimes y)^{t}=(xy^{t})^{t}=yx^{t}=y\otimes x$
for $x,y\in\mathbb{O}$. Note that a notational ambiguity could arise from this identification, since now we have defined two different binary operations on $\mathbb{O}\otimes\mathbb{O}=\mathbb{R}(8)$, oct-octonionic multiplication and the usual matrix multiplication. We avoid this ambiguity by writing $x\ast y$ for oct-octonionic multiplication, where $x,y\in\mathbb{O}\otimes\mathbb{O}$ and by denoting the matrix product by juxtaposition.
For $A=(P,Q)\in\operatorname{\operatorname{\mathfrak{so}}}(8)\oplus\operatorname{\operatorname{\mathfrak{so}}}(8)$ and $X\in\mathbb{O}\otimes\mathbb{O}=\mathbb{R}(8)$, we define
$A.X:=PX-XQ=PX+XQ^{t},$ (3.2)
where the terms on the right hand side are defined by using the usual matrix multiplication, viewing $P,Q,X$ as real $8\times 8$ matrices. In particular, the Lie algebra $\operatorname{\operatorname{\mathfrak{so}}}(8)\oplus\operatorname{\operatorname{\mathfrak{so}}}(8)$ acts on $\mathbb{R}^{8}\otimes\mathbb{R}^{8}$ by the outer tensor product representation of the two standard representations of the two summands in $\operatorname{\operatorname{\mathfrak{so}}}(8)\oplus\operatorname{\operatorname{\mathfrak{so}}}(8)$, i.e. we have
$(P,Q).(r\otimes s)=(P,Q).sr^{t}=Prs^{t}+rs^{t}Q^{t}=(Pr)\otimes s+r\otimes(Qs)$ (3.3)
for $(P,Q)\in\operatorname{\operatorname{\mathfrak{so}}}(8)\times\operatorname{\operatorname{\mathfrak{so}}}(8)$ $X=r\otimes s$, $r,s\in\mathbb{O}$, We furthermore define for $X,Y\in\mathbb{R}(8)$:
$X\curlywedge Y:=(XY^{t}-YX^{t},X^{t}Y-Y^{t}X)\in\operatorname{\operatorname{\mathfrak{so}}}(8)\times\operatorname{\operatorname{\mathfrak{so}}}(8).$ (3.4)
Note that in this definition, matrix multiplication is understood (instead of oct-octonionic multiplication). Let $p,q,r,s\in\mathbb{O}$. We compute for later use:
$(p\otimes q)$ $\curlywedge(r\otimes s)=$
$=((p\otimes q)(r\otimes s)^{t}-(r\otimes s)(p\otimes q)^{t},(p\otimes q)^{t}(r\otimes s)-(r\otimes s)^{t}(p\otimes q))=$
$=(pq^{t}sr^{t}-rs^{t}qp^{t},qp^{t}rs^{t}-sr^{t}pq^{t})=$
$=(\langle q,s\rangle p\wedge r,\langle p,r\rangle q\wedge s).$ (3.5)
Lemma 3.1.
Let $(\operatorname{\operatorname{\mathfrak{so}}}(8)\oplus\operatorname{\operatorname{\mathfrak{so}}}(8))\times\mathbb{R}(8)$ be equipped with the binary operation defined by
$[(A,X),(B,Y)]:=(AB-BA-4X\curlywedge Y,A.Y-B.X),$
for $A,B\in\operatorname{\operatorname{\mathfrak{so}}}(8)\oplus\operatorname{\operatorname{\mathfrak{so}}}(8)$ and $X,Y\in\mathbb{R}(8)$. Then the real algebra defined in this fashion is isomorphic to the Lie algebra $\operatorname{\operatorname{\mathfrak{so}}}(16)$.
Proof.
We will check that an isomorphism is given by the map
$((P,Q),X)\mapsto\begin{pmatrix}P&2X\\ -2X^{t}&Q\\ \end{pmatrix}.$
Indeed, let us compute the bracket
$\left[\begin{pmatrix}P&2X\\ -2X^{t}&Q\\ \end{pmatrix}\right.,\left.\begin{pmatrix}R&2Y\\ -2Y^{t}&S\\ \end{pmatrix}\right]=$
$\;\;\;\;=\begin{pmatrix}PR-RP-4XY^{t}+4YX^{t}&2PY-2YQ-2RX+2XS\\ 2Y^{t}P-2QY^{t}-2X^{t}R+2SX^{t}&QS-SQ-4X^{t}Y+4Y^{t}X\\ \end{pmatrix}=$
$\;\;\;\;=\begin{pmatrix}[P,R]-4(XY^{t}-YX^{t})&2(P,Q).Y-2(R,S).X\\ -(2(P,Q).Y-2(R,S).X)^{t}&[Q,S]-4(X^{t}Y-Y^{t}X)\\ \end{pmatrix}$
This proves the statement of the lemma ∎
We define conjugation $\mathbb{O}\otimes\mathbb{O}\to\mathbb{O}\otimes\mathbb{O}$, $x\mapsto\bar{x}$, of oct-octonions to be the $\mathbb{R}$-linear map given by
$\overline{a\otimes b}:=\bar{a}\otimes\bar{b}$
for $a,b\in\mathbb{O}$. Conjugation of oct-octonions is an anti-automorphism of $\mathbb{O}\otimes\mathbb{O}$, i.e. we have $\overline{xy}=\bar{y}\bar{x}$ for $x,y\in\mathbb{O}\otimes\mathbb{O}$, this follows from the fact that conjugation of octonions is an antiautomorphism of $\mathbb{O}$. We define an outer automorphism of order three of $\operatorname{\operatorname{\mathfrak{so}}}(8)\oplus\operatorname{\operatorname{\mathfrak{so}}}(8)$ by
$\Lambda\colon\operatorname{\operatorname{\mathfrak{so}}}(8)\oplus\operatorname{\operatorname{\mathfrak{so}}}(8)$ $\to\operatorname{\operatorname{\mathfrak{so}}}(8)\oplus\operatorname{\operatorname{\mathfrak{so}}}(8),$ (3.6)
$(P,Q)$ $\mapsto(\lambda(P),\lambda(Q)).$
4. The Lie algebra $\mathfrak{e}_{8}$
Theorem 4.1.
The binary operation on $\mathscr{A}=(\operatorname{\operatorname{\mathfrak{so}}}(8)\oplus\operatorname{\operatorname{\mathfrak{so}}}(8))\times(\mathbb{O}\otimes\mathbb{O})^{3}$ defined by
$[(A,u,v,w),(B,x,y,z)]=(C,r,s,t)$ (4.1)
where
$C$ $=[A,B]-4u\curlywedge x-4\Lambda^{2}(v\curlywedge y)-4\Lambda(w\curlywedge z),$
$r$ $=A.x-B.u+\overline{v\ast z}-\overline{y\ast w},$
$s$ $=\Lambda(A).y-\Lambda(B).v+\overline{w\ast x}-\overline{z\ast u},$
$t$ $=\Lambda^{2}(A).z-\Lambda^{2}(B).w+\overline{u\ast y}-\overline{x\ast v},$
and where $A,B\in\operatorname{\operatorname{\mathfrak{so}}}(8)\oplus\operatorname{\operatorname{\mathfrak{so}}}(8)$, $u,v,w,x,y,z\in\mathbb{O}\otimes\mathbb{O}$, is $\mathbb{R}$-bilinear, skew symmetric and satisfies the Jacobi identity. The real Lie algebra defined in this way is isomorphic to the Lie algebra of the compact exceptional simple Lie group of type $\rm E_{8}$.
It follows directly by inspection that the binary operation (4.1) is $\mathbb{R}$-bilinear and skew symmetric. To prove Theorem 4.1 we will verify that the Jacobi identity holds and that the resulting Lie algebra is a simple Lie algebra of dimension $248$ which carries an $\operatorname{ad}$-invariant inner product. Until we have completed the proof, we will write $\mathscr{A}$ for the real algebra defined by the operation (4.1).
Lemma 4.2.
The linear map $\tau\colon\mathscr{A}\to\mathscr{A}$, given by
$\tau(A,x,y,z)=(\Lambda(A),y,z,x)$
is an automorphism of the real algebra $\mathscr{A}$.
Proof.
Follows directly by inspection of the formula (4.1) and using the fact that $\Lambda$ is an automorphism of order three of $\operatorname{\operatorname{\mathfrak{so}}}(8)\oplus\operatorname{\operatorname{\mathfrak{so}}}(8)$. ∎
We define an action of the Lie group $\operatorname{Spin}(8)\times\operatorname{Spin}(8)$ on $\mathbb{R}(8)=\mathbb{O}\otimes\mathbb{O}$ by
$\Theta(x):=(\theta_{1},\theta_{2})(u\otimes v):=\theta_{1}(u)\otimes\theta_{2}(v)$
for $\Theta=(\theta_{1},\theta_{2})\in\operatorname{Spin}(8)\times\operatorname{Spin}(8)$ and $x=u\otimes v\in\mathbb{R}(8)$ and where it is understood that an element of $\operatorname{Spin}(8)$ acts by the standard representation of $\operatorname{Spin}(8)$ on $\mathbb{R}^{8}$. Since $\operatorname{Spin}(8)\times\operatorname{Spin}(8)$ is simply connected, its automorphism group is canonically isomorphic to the automorphism group of $\operatorname{\operatorname{\mathfrak{so}}}(8)\oplus\operatorname{\operatorname{\mathfrak{so}}}(8)$ and we denote the automorphism of $\operatorname{Spin}(8)\times\operatorname{Spin}(8)$ given by $\Lambda\in\operatorname{Aut}(\operatorname{\operatorname{\mathfrak{so}}}(8)\oplus\operatorname{\operatorname{\mathfrak{so}}}(8))$ also by $\Lambda$. Define an action of $\operatorname{Spin}(8)\times\operatorname{Spin}(8)$ on $\mathscr{A}$ by
$\Theta(A,x,y,z)=(\operatorname{Ad}_{\Theta}(A),\Theta(x),\Lambda(\Theta)(y),\Lambda^{2}(\Theta)(z))$
for $\Theta\in\operatorname{Spin}(8)\times\operatorname{Spin}(8)$.
Lemma 4.3.
We have
$\tau(\Theta(A,x,y,z))=\Lambda(\Theta)(\tau(A,x,y,z))$
for all $\Theta\in\operatorname{Spin}(8)\times\operatorname{Spin}(8)$, $(A,x,y,z)\in\mathscr{A}$.
Proof.
We compute
$\tau(\Theta(A,x,y,z))$ $=\tau(\operatorname{Ad}_{\Theta}(A),\Theta(x),\Lambda(\Theta)(y),\Lambda^{2}(\Theta)(z))=$
$=(\operatorname{Ad}_{\Lambda(\Theta)}(\Lambda(A)),\Lambda(\Theta)(y),\Lambda^{2}(\Theta)(z),\Theta(x))=$
$=\Lambda(\Theta)(\Lambda(A),y,z,x)=\Lambda(\Theta)(\tau(A,x,y,z)).\qed$
Lemma 4.4.
The map
$(A,x,y,z)\mapsto\Theta(A,x,y,z)$
is an automorphism of the real algebra $\mathscr{A}$ for any $\Theta\in\operatorname{Spin}(8)\times\operatorname{Spin}(8)$.
Proof.
Let $\Theta=(\theta_{1},\theta_{2})\in\operatorname{Spin}(8)\times\operatorname{Spin}(8)$. We first prove the automorphism property in some special cases, then we use linearity and the map $\tau$ to deduce the statement of the lemma. Note that
$[\Theta(A,0,0,0),\Theta(B,0,0,0)]=[(\operatorname{Ad}_{\Theta}(A),0,0,0),(\operatorname{Ad}_{\Theta}(B),0,0,0)]=$
$=(\operatorname{Ad}_{\Theta}([A,B]),0,0,0)=\Theta[(A,0,0,0),(B,0,0,0)].$
We compute
$[\Theta$ $(A,0,0,0),\Theta(0,x,y,z)]=[(\operatorname{Ad}_{\Theta}(A),0,0,0),(0,\Theta(x),\Lambda(\Theta)(y),\Lambda^{2}(\Theta)(z))]=$
$=(0,\operatorname{Ad}_{\Theta}(A)(\Theta(x)),\Lambda(\operatorname{Ad}_{\Theta}(A))(\Lambda(\Theta)(y)),\Lambda^{2}(\operatorname{Ad}_{\Theta}(A))(\Lambda^{2}(\Theta)(z)))=$
$=(0,\Theta(Ax),\Lambda(\Theta)(\Lambda(A)y),\Lambda^{2}(\Theta)(\Lambda^{2}(A)z))=$
$=\Theta(0,Ax,\Lambda(A)y,\Lambda^{2}(A)z)=\Theta[(A,0,0,0),(0,x,y,z)].$
It follows from the principle of triality, see [12, §2-3], that
$\theta(a)\lambda(\theta)(b)=\kappa\circ\lambda^{2}(\theta)(ab)$ (4.2)
for all $a,b\in\mathbb{O}$ and all $\theta\in\operatorname{Spin}(8)$. We will use this in the following computation. Let $u=p\otimes q$ and $y=r\otimes s$, where $p,q,r,s\in\mathbb{O}$.
$[\Theta$ $(0,u,0,0),\Theta(0,0,y,0)]=[(0,\Theta(u),0,0),(0,0,\Lambda(\Theta)(y),0)]=$
$=(0,0,0,\overline{\Theta(u)\ast\Lambda(\Theta)(y)})=$
$=(0,0,0,\overline{(\theta_{1}(p)\otimes\theta_{2}(q))\ast(\lambda(\theta_{1})(r)\otimes\lambda(\theta_{2})(s))})=$
$=(0,0,0,\overline{\theta_{1}(p)\lambda(\theta_{1})(r)\otimes\theta_{2}(q)\lambda(\theta_{2})(s)})=$
$=(0,0,0,\overline{\kappa\circ\lambda^{2}(\theta_{1})(pr)\otimes\kappa\circ\lambda^{2}(\theta_{2})(qs)})=$
$=(0,0,0,\overline{\kappa\circ\lambda^{2}(\theta_{1})(pr)}\otimes\overline{\kappa\circ\lambda^{2}(\theta_{2})(qs)})=$
$=(0,0,0,\lambda^{2}(\theta_{1})(\overline{pr})\otimes\lambda^{2}(\theta_{2})(\overline{qs}))=$
$=(0,0,0,\Lambda^{2}(\Theta)(\overline{u\ast y}))=\Theta(0,0,0,\overline{u\ast y})=\Theta[(0,u,0,0),(0,0,y,0)],$
where we have used that $\kappa(\theta)(x)=\overline{\theta(\bar{x})}$ for all $\theta\in\operatorname{Spin}(8),x\in\mathbb{O}$. Finally, assuming $u=p\otimes q$, $x=r\otimes s$ for $p,q,r,s\in\mathbb{O}$ and $\Theta=(\theta_{1},\theta_{2})$ we obtain
$[\Theta$ $(0,u,0,0),\Theta(0,x,0,0)]=[(0,\Theta(u),0,0),(0,\Theta(x),0,0)]=$
$=[(0,\Theta(u),0,0),(0,\Theta(x),0,0)]=$
$=(-4\Theta(u)\curlywedge\Theta(x),0,0,0)=$
$=(-4(\theta_{1}pq^{t}\theta_{2}^{t})\curlywedge(\theta_{1}rs^{t}\theta_{2}^{t}),0,0,0)=$
$=(-4(\theta_{1}pq^{t}\theta_{2}^{t}\theta_{2}sr^{t}\theta_{1}^{t}-\theta_{1}rs^{t}\theta_{2}^{t}\theta_{2}qp^{t}\theta_{1}^{t},\theta_{2}qp^{t}\theta_{1}^{t}\theta_{1}rs^{t}\theta_{2}^{t}-\theta_{2}sr^{t}\theta_{1}^{t}\theta_{1}pq^{t}\theta_{2}^{t}),0,0,0)=$
$=(-4(\theta_{1}pq^{t}sr^{t}\theta_{1}^{t}-\theta_{1}rs^{t}qp^{t}\theta_{1}^{t},\theta_{2}qp^{t}rs^{t}\theta_{2}^{t}-\theta_{2}sr^{t}pq^{t}\theta_{2}^{t}),0,0,0)=$
$=(-4\operatorname{Ad}_{\Theta}(pq^{t}sr^{t}-rs^{t}qp^{t},qp^{t}rs^{t}-sr^{t}pq^{t}),0,0,0)=$
$=(-4\operatorname{Ad}_{\Theta}(u\curlywedge x),0,0,0)=\Theta([(0,u,0,0),(0,x,0,0)]).$
Using the skew symmetry and bilinearity of the bracket, the fact that $\tau$ is an automorphism of $\mathscr{A}$ and Lemma 4.3, the statement of the lemma now follows. ∎
Lemma 4.5.
The bracket operation on $\mathscr{A}$ satisfies the Jacobi identity, i.e. we have
$[\xi,[\eta,\zeta]]+[\eta,[\zeta,\xi]]+[\zeta,[\xi,\eta]]=0$ (4.3)
for all $\xi,\eta,\zeta\in\mathscr{A}$.
Proof.
Since the left-hand side of the Jacobi identity (4.3) is trilinear in $\xi,\eta,\zeta$, it suffices to show that it holds for all triples of vectors $(\xi,\eta,\zeta)$ where each one of the vectors $\xi,\eta,\zeta$ is an element from one of the four factors of $(\operatorname{\operatorname{\mathfrak{so}}}(8)\oplus\operatorname{\operatorname{\mathfrak{so}}}(8))\times(\mathbb{O}\otimes\mathbb{O})\times(\mathbb{O}\otimes\mathbb{O})\times(\mathbb{O}\otimes\mathbb{O})$. Since the map $\tau$ is an automorphism of $\mathscr{A}$, we may furthermore assume the three vectors $\xi,\eta,\zeta$ are either from $(\operatorname{\operatorname{\mathfrak{so}}}(8)\oplus\operatorname{\operatorname{\mathfrak{so}}}(8))\times(\mathbb{O}\otimes\mathbb{O})\times(\mathbb{O}\otimes\mathbb{O})\times\{0\}$ or from $\{0\}\times(\mathbb{O}\otimes\mathbb{O})\times(\mathbb{O}\otimes\mathbb{O})\times(\mathbb{O}\otimes\mathbb{O})$.
  1. (i)
    Assume first $\xi,\eta,\zeta\in(\operatorname{\operatorname{\mathfrak{so}}}(8)\oplus\operatorname{\operatorname{\mathfrak{so}}}(8))\times(\mathbb{O}\otimes\mathbb{O})\times(\mathbb{O}\otimes\mathbb{O})\times\{0\}$. In the special cases where the vectors $\xi,\eta,\zeta$ are only taken from the first two factors or from the first and third factor, it follows from Lemma 3.1 (using the fact that $\tau$ is an automorphism of $\mathscr{A}$ in the second case) that 4.3 holds. Therefore, we assume
    $\xi=(A,0,0,0),\quad\eta=(0,x,0,0),\quad\zeta=(0,0,y,0).$
    We compute the three summands on the left-hand side of (4.3):
    $[(A,0,0,0),[(0,x,0,0),(0,0,y,0)]]$ $=(0,0,0,\Lambda^{2}(A).\overline{x\ast y}),$
    $[(0,x,0,0),[(0,0,y,0),(A,0,0,0)]]$ $=(0,0,0,-\overline{x\ast(\Lambda(A).y)}),$
    $[(0,0,y,0),[(A,0,0,0),(0,x,0,0)]]$ $=(0,0,0,-\overline{(A.x)\ast y}).$
    Summing up the fourth components of these elements, and writing $A=(P,Q)$, $x=p\otimes q$, $y=r\otimes s$, where $P,Q\in\operatorname{\operatorname{\mathfrak{so}}}(8)$, $p,q,r,s\in\mathbb{O}$, we get
    $\Lambda^{2}$ $(A).\overline{x\ast y}-\overline{x\ast(\Lambda(A).y)}-\overline{(A.x)\ast y}=$
    $=\Lambda^{2}(A).\overline{pr\otimes qs}-\overline{x\ast(\lambda(P)y-y\lambda(Q))}-\overline{(Px-xQ)\ast y}=$
    $=\lambda^{2}(P)\overline{pr}\otimes\overline{qs}+\overline{pr}\otimes\lambda^{2}(Q)\overline{qs}-\overline{(p\otimes q)\ast(\lambda(P)r\otimes s+r\otimes\lambda(Q)s)}-$
    $\;\;-\overline{(Pp\otimes q+p\otimes Qq)\ast(r\otimes s)}=$
    $=\lambda^{2}(P)\overline{pr}\otimes\overline{qs}+\overline{pr}\otimes\lambda^{2}(Q)\overline{qs}-\overline{p\lambda(P)r\otimes qs+pr\otimes q\lambda(Q)s}-$
    $\ -\overline{(Pp)r\otimes qs+pr\otimes(Qq)s)}=$
    $=\overline{(\kappa\circ\lambda^{2}(P)(pr)-p\lambda(P)r-(Pp)r)\otimes qs}+$
    $\;\;+\overline{pr\otimes(\kappa\circ\lambda^{2}(Q)(qs)-q\lambda(Q)s-(Qq)s)}=$
    $=\overline{0\otimes qs}+\overline{pr\otimes 0}=0,$
    where we have used the infinitesimal principle of triality, see [12, §2, Thm. 1], which implies $(Au)v+u\lambda(A)v=\kappa\circ\lambda^{2}(A)(uv)$ for all $A\in\operatorname{\operatorname{\mathfrak{so}}}(8)$ and all $u,v\in\mathbb{O}$.
  2. (ii)
    It now remains to study the case where the three elements $\xi,\eta,\zeta$ from $\{0\}\times(\mathbb{O}\otimes\mathbb{O})\times(\mathbb{O}\otimes\mathbb{O})\times(\mathbb{O}\otimes\mathbb{O})$. Again by using the automorphism $\tau$, it suffices to study the two subcases where three elements are from three distinct factors in $(\mathbb{O}\otimes\mathbb{O})\times(\mathbb{O}\otimes\mathbb{O})\times(\mathbb{O}\otimes\mathbb{O})$ or from only two distinct factors.
    1. (a)
      Let us assume
      $\xi=(0,x,0,0),\quad\eta=(0,0,y,0),\quad\zeta=(0,0,0,z)$
      where $x=e_{i}\otimes e_{j}$, $y=e_{k}\otimes e_{\ell}$, $z=t\otimes u$, $t,u\in\mathbb{O}$. The group $\operatorname{Spin}(8)\times\operatorname{Spin}(8)$ acts as a group of automorphisms on $\mathscr{A}$ by Lemma 4.4 and we may use this action to assume $x=e\otimes e$ and $y=e\otimes e$, since $\operatorname{Spin}(8)$ acts transitively on the product of unit spheres $\mathrm{S}^{7}\times\mathrm{S}^{7}\subset\mathbb{R}^{8}\oplus\mathbb{R}^{8}$ by the sum of any two of its inequivalent irreducible 8-dimensional representations. We compute, using (3.5),
      $[(0,x,0,0)$ $,[(0,0,y,0),(0,0,0,z)]]=$
      $=[(0,e\otimes e,0,0),(0,\overline{t\otimes u},0,0)]=$
      $=(-4(e\otimes e)\curlywedge(\bar{t}\otimes\bar{u}),0,0,0)=$
      $=(-4(\langle e,\bar{u}\rangle e\wedge\bar{t},\langle e,\bar{t}\rangle e\wedge\bar{u}),0,0,0),$
      $[(0,0,y,0)$ $,[(0,0,0,z),(0,x,0,0)]]=$
      $=[(0,0,e\otimes e,0),(0,0,\overline{t\otimes u},0)]=$
      $=(-4\Lambda^{2}((e\otimes e)\curlywedge(\bar{t}\otimes\bar{u})),0,0,0),$
      $=(-4\Lambda^{2}((\langle e,\bar{u}\rangle\bar{t}\wedge e,\langle e,\bar{t}\rangle\bar{u}\wedge e)),0,0,0)=$
      $=(-4(\langle e,\bar{u}\rangle\lambda^{2}(\bar{t}\wedge e),\langle e,\bar{t}\rangle\lambda^{2}(\bar{u}\wedge e)),0,0,0),$
      $[(0,0,0,z)$ $,[(0,x,0,0),(0,0,y,0)]]=$
      $=[(0,0,0,t\otimes u),(0,0,0,e\otimes e)]=$
      $=(-4\Lambda((t\otimes u)\curlywedge(e\otimes e)),0,0,0)=$
      $=(-4\Lambda(\langle e,\bar{u}\rangle\bar{t}\wedge e,\langle e,\bar{t}\rangle\bar{u}\wedge e)),0,0,0)=$
      $=(-4(\langle e,\bar{u}\rangle\lambda(\bar{t}\wedge e),\langle e,\bar{t}\rangle\lambda(\bar{u}\wedge e))),0,0,0).$
      The sum of these three elements is zero if
      00 0 $=e\wedge\bar{t}+\lambda^{2}(\bar{t}\wedge e)+\lambda(\bar{t}\wedge e)=e\wedge\bar{u}+\lambda^{2}(\bar{u}\wedge e)+\lambda(\bar{u}\wedge e).$
      Note that, by linearity, we may assume $t,u\in\operatorname{Pu}(\mathbb{O})$, since $\bar{t}\wedge e=0$ if $t\in\mathbb{R}$ and $\bar{u}\wedge e=0$ if $u\in\mathbb{R}$. Then the above two equations are equivalent to
      $0=e\wedge 2t+R_{t}+L_{t}=e\wedge 2u+R_{u}+L_{u}.$
      It can be directly verified that $2t\wedge e=R_{t}+L_{t}$ holds for $t\in\operatorname{Pu}(\mathbb{O})$, see [11, Equation (2.2)].
    2. (b)
      Now consider
      $\xi=(0,x,0,0),\quad\eta=(0,y,0,0),\quad\zeta=(0,0,z,0).$
      We compute
      $[(0,x,0,0)$ $,[(0,y,0,0),(0,0,z,0)]]=$
      $=[(0,x,0,0),(0,0,0,\overline{y\ast z})]=(0,0,-\bar{x}\ast(y\ast z),0),$
      $[(0,y,0,0)$ $,[(0,0,z,0),(0,x,0,0)]]=$
      $=[(0,y,0,0),(0,0,0,-\overline{x\ast z})]=(0,0,\bar{y}\ast(x\ast z),0),$
      $[(0,0,z,0)$ $,[(0,x,0,0),(0,y,0,0)]]=$
      $=[(0,0,z,0),(-4x\curlywedge y,0,0,0)]=(0,0,4\Lambda(x\curlywedge y).z,0).$
      Let $y=p\otimes q$. Using the $\operatorname{Spin}(8){\times}\operatorname{Spin}(8)$-action, we may assume $x=z=e\otimes e$. Then the sum of the third components of the above three elements becomes
      $-y$ $+\bar{y}+4\Lambda((e\otimes e)\curlywedge(p\otimes q)).(e\otimes e)=$
      $=-y+\bar{y}+4(\langle e,q\rangle\lambda(e\wedge p),\langle e,p\rangle\lambda(e\wedge q)).(e\otimes e)=$
      $=-y+\bar{y}+4(\langle e,q\rangle\lambda(e\wedge p)e\otimes e+\langle e,p\rangle e\otimes\lambda(e\wedge q)e)=$
      $=-y+\bar{y}-4(\langle e,q\rangle\tfrac{1}{2}L_{\operatorname{Pu}(p)}(e)\otimes e+\langle e,p\rangle e\otimes\tfrac{1}{2}L_{\operatorname{Pu}(q)}(e))=$
      $=-p\otimes q+\bar{p}\otimes\bar{q}-2(\langle e,q\rangle\operatorname{Pu}(p)\otimes e+\langle e,p\rangle e\otimes\operatorname{Pu}(q))=0.$
      To see that the term in the last line is zero, one may use the identification (3.1) to write the summands as matrices. Then $-p\otimes q+\bar{p}\otimes\bar{q}$ is a real $8\times 8$ matrix which has non-zero entries only in the first row and the first column and so is $2(\langle e,q\rangle\operatorname{Pu}(p)\otimes e+\langle e,p\rangle e\otimes\operatorname{Pu}(q))$; it is easy to see that these two matrices are equal.
    3. (c)
      Finally, we compute
      $[(0,x,0,0)$ $,[(0,0,y,0),(0,0,z,0)]]=$
      $=[(0,x,0,0),(-4\Lambda^{2}(y\curlywedge z),0,0,0)]=(0,4\Lambda^{2}(y\curlywedge z).x,0,0),$
      $[(0,0,y,0)$ $,[(0,0,z,0),(0,x,0,0)]]=$
      $=[(0,0,y,0),(0,0,0,-\overline{x\ast z})]=(0,-(x\ast z)\ast\bar{y},0,0),$
      $[(0,0,z,0)$ $,[(0,x,0,0),(0,0,y,0)]]=$
      $=[(0,0,z,0),(0,0,0,\overline{x\ast y})]=(0,(x\ast y)\ast\bar{z},0,0)$
      We may again assume $x=z=e\otimes e$, $y=p\otimes q$. The sum of the second components of above three elements of $\mathscr{A}$ is then
      44 44 $\Lambda^{2}(y\curlywedge z).x-(x\ast z)\ast\bar{y}+(x\ast y)\ast\bar{z}=$
      $=4\Lambda^{2}((p\otimes q)\curlywedge(e\otimes e)).(e\otimes e)-\bar{p}\otimes\bar{q}+p\otimes q=$
      $=4\Lambda^{2}(\langle q,e\rangle p\wedge e,\langle p,e\rangle q\wedge e).(e\otimes e)-\bar{p}\otimes\bar{q}+p\otimes q=$
      $=4(\langle q,e\rangle\lambda^{2}(p\wedge e),\langle p,e\rangle\lambda^{2}(q\wedge e)).(e\otimes e)-\bar{p}\otimes\bar{q}+p\otimes q=$
      $=4\langle q,e\rangle\lambda^{2}(p\wedge e)(e)\otimes e+4\langle p,e\rangle e\otimes\lambda^{2}(q\wedge e)(e)-\bar{p}\otimes\bar{q}-+p\otimes q=$
      $=2\langle q,e\rangle R_{\operatorname{Pu}(p)}(e)\otimes e+2\langle p,e\rangle e\otimes R_{\operatorname{Pu}(q)}(e)-\bar{p}\otimes\bar{q}+p\otimes q=$
      $=2\langle q,e\rangle\operatorname{Pu}(p)\otimes e+2\langle p,e\rangle e\otimes\operatorname{Pu}(q)-\bar{p}\otimes\bar{q}+p\otimes q=0,$
      as above.
We have shown that the Jacobi identity holds for the real algebra $\mathscr{A}$. ∎
We define a scalar product on the Lie algebra $\mathscr{A}$ by
$\langle((A,B),u,v,w),((C,D),x,y,z)\rangle=8\operatorname{tr}(ux^{t}+vy^{t}+wz^{t})-\operatorname{tr}(AC)-\operatorname{tr}(BD).$ (4.4)
This scalar product is easily seen to be invariant under $\tau$ and the $\operatorname{Spin}(8)\times\operatorname{Spin}(8)$-action.
Lemma 4.6.
The scalar product (4.4) on the Lie algebra $\mathscr{A}$ is $\operatorname{ad}$-invariant.
Proof.
Let $((A,B),u,v,w),((C,D),x,y,z)\in\mathscr{A}$ and let $P,Q\in\operatorname{\operatorname{\mathfrak{so}}}(8)$. We compute
$\langle[((A,B),u,v,w),((P,Q),0,0,0)],((C,D),x,y,z)\rangle=$
$=\langle([(A,B),(P,Q)],-(P,Q).u,-\Lambda(P,Q).v,-\Lambda^{2}(P,Q).w),((C,D),x,y,z)\rangle=$
$=8\operatorname{tr}((-Pu+uQ)x^{t}+(-\lambda(P)v+v\lambda(Q))y^{t}+(-\lambda^{2}(P)w+w\lambda^{2}(Q))z^{t})-$
$\;\;-\operatorname{tr}(APC-PAC)-\operatorname{tr}(BQD-QBD)=$
$=8\operatorname{tr}(u(Px-Qx)^{t}+v(\lambda(P)y-y\lambda(Q))^{t}+w(\lambda^{2}(P)z-z\lambda^{2}(Q))^{t})-$
$\;\;-\operatorname{tr}(APC-ACP)-\operatorname{tr}(BQD-BDQ)=$
$=\langle((A,B),u,v,w),([(P,Q),(C,D)],(P,Q).x,\Lambda(P,Q).y,\Lambda^{2}(P,Q).z)\rangle=$
$=\langle((A,B),u,v,w),[((P,Q),0,0,0),((C,D),x,y,z)]\rangle.$
Furthermore, we have
$\langle[((A,B),u,v,w),(0,e\otimes e,0,0)],((C,D),x,y,z)\rangle=$
$=\langle(4(e\otimes e)\curlywedge u,(A,B).(e\otimes e),\bar{w},-\bar{v}),((C,D),x,y,z)\rangle=$
$=\langle((4ee^{t}u^{t}-4uee^{t},4ee^{t}u-u^{t}ee^{t}),Aee^{t}+ee^{t}B^{t},\bar{w},-\bar{v}),((C,D),x,y,z)\rangle=$
$=8\operatorname{tr}(Aee^{t}x^{t}+ee^{t}B^{t}x^{t}+\bar{w}y^{t}-\bar{v}z^{t})-$
$\;\;-\operatorname{tr}(4ee^{t}u^{t}C-4uee^{t}C)-\operatorname{tr}(4ee^{t}uD-4u^{t}ee^{t}D)=$
$=\operatorname{tr}(8Aee^{t}x^{t}+8B^{t}x^{t}ee^{t}+8Cuee^{t}+8Du^{t}ee^{t}-8\bar{v}z^{t}+8\bar{w}y^{t})=$
$=8\operatorname{tr}(-uee^{t}C^{t}-uDee^{t}-v\bar{z}^{t}+w\bar{y}^{t})-$
$\;\;-\operatorname{tr}(4Axee^{t}-4Aee^{t}x^{t})-\operatorname{tr}(4Bx^{t}ee^{t}-4Bee^{t}x)=$
$=\langle((A,B),u,v,w),((4xee^{t}-4ee^{t}x^{t},4x^{t}ee^{t}-4ee^{t}x),-Cee^{t}-ee^{t}D^{t},-\bar{z},\bar{y})\rangle=$
$=\langle((A,B),u,v,w),(4x\curlywedge(e\otimes e),-(C,D).(e\otimes e),-\bar{z},\bar{y})\rangle=$
$=\langle((A,B),u,v,w),[(0,e\otimes e,0,0),((C,D),x,y,z)]\rangle.$
Here we have used the fact that $\operatorname{tr}(\bar{x}y^{t})=\operatorname{tr}(x\bar{y}^{t})$ for $x,y\in\mathbb{O}\otimes\mathbb{O}$. Since the scalar product on $\mathscr{A}$ is invariant under $\tau$ and under the $\operatorname{Spin}(8){\times}\operatorname{Spin}(8)$-action, the above calculations suffice to prove the $\operatorname{ad}$-invariance of the scalar product. ∎
Proof of Theorem 4.1.
We have shown in Lemma 4.5 that the skew-symmetric operation of the real algebra $\mathscr{A}$ satisfies the Jacobi identity. Therefore, $\mathscr{A}$ with the bracket defined in (4.1) is a real Lie algebra. It can be seen from the bracket formula (4.1) that if the adjoint representation of $\mathscr{A}$ is restricted to the subalgebra $(\operatorname{\operatorname{\mathfrak{so}}}(8)\oplus\operatorname{\operatorname{\mathfrak{so}}}(8))\times\{0\}\times\{0\}\times\{0\}$, then the resulting representation on $\mathscr{A}$ is a direct sum of $5$ inequivalent irreducible modules. Thus any non-trivial ideal of $\mathscr{A}$ is a direct sum of one or more of these $5$ summands. However, inspection of the formula (4.1) shows that the only such ideal is $\mathscr{A}$ itself and thus the Lie algebra $\mathscr{A}$ is simple. We have $\dim(\mathscr{A})=248$; using the formulae $\dim(\mathrm{A}_{n})=n(n+2)$, $\dim(\mathrm{B}_{n})=\dim(\mathrm{C}_{n})=n(2n+1)$, $\dim(\mathrm{D}_{n})=n(2n-1)$, $\dim(\mathrm{G}_{2})=14$, $\dim(\mathrm{F}_{4})=52$, $\dim(\mathrm{E}_{6})=78$, $\dim(\mathrm{E}_{7})=133$, $\dim(\mathrm{E}_{8})=248$, it is easy to see that any simple real Lie algebra of dimension $248$ is of type $\mathrm{E}_{8}$. There are three real forms of the complex simple Lie algebra of type $\mathrm{E}_{8}$, the non-compact real forms $\operatorname{\mathfrak{e}}_{8(8)}$ and $\operatorname{\mathfrak{e}}_{8(-24)}$ and the compact form $\operatorname{\mathfrak{e}}_{8(-248)}$. Since we have shown in Lemma 4.6 that there exists an $\operatorname{ad}$-invariant scalar product on $\mathscr{A}$, it follows that $\mathscr{A}$ is isomorphic to the compact real form $\operatorname{\mathfrak{e}}_{8(-248)}$. ∎
5. The noncompact simple Lie algebra $\mathfrak{e}_{8(8)}$
Using the duality of Riemannian symmetric spaces, we immediately obtain the bracket formula for the real noncompact simple Lie algebra $\mathfrak{e}_{8(8)}$.
Corollary 5.1.
The binary operation on $(\operatorname{\operatorname{\mathfrak{so}}}(8)\oplus\operatorname{\operatorname{\mathfrak{so}}}(8))\times(\mathbb{O}\otimes\mathbb{O})^{3}$ defined by
$[(A,u,v,w),(B,x,y,z)]=(C^{*},r^{*},s,t)$ (5.1)
where
$C^{*}$ $=[A,B]-4u\curlywedge x+4\Lambda^{2}(v\curlywedge y)+4\Lambda(w\curlywedge z),$
$r^{*}$ $=A.x-B.u-\overline{v\ast z}+\overline{y\ast w},$
$s$ $=\Lambda(A).y-\Lambda(B).v+\overline{w\ast x}-\overline{z\ast u},$
$t$ $=\Lambda^{2}(A).z-\Lambda^{2}(B).w+\overline{u\ast y}-\overline{x\ast v},$
where $A,B\in\operatorname{\operatorname{\mathfrak{so}}}(8)\oplus\operatorname{\operatorname{\mathfrak{so}}}(8)$, $u,v,w,x,y,z\in\mathbb{O}\otimes\mathbb{O}$, defines the bracket operation of a real Lie algebra isomorphic to the noncompact exceptional simple algebra $\mathfrak{e}_{8(8)}$.
Proof.
Let $\mathfrak{g}=\mathfrak{e}_{8}$, the compact form. It is easy to see from the bracket formula (4.1) and Lemma 3.1 that the map $(A,u,v,w)\mapsto(A,u,-v,-w)$ is an involutive automorphism of $\mathfrak{g}$, corresponding to the Riemannian symmetric space ${\rm E}_{8}/\operatorname{Spin}(16)$. Let $\mathfrak{g}=\mathfrak{k}\oplus\mathfrak{p}$ be the decomposition into the $(+1)$-eigenspace and $(-1)$-eigenspace. It follows that $\mathfrak{g}=\mathfrak{k}\oplus i\mathfrak{p}\subset\mathfrak{g}\otimes\mathbb{C}$ is the Lie algebra of the isometry group of the dual symmetric space ${\rm E}_{8(8)}/\operatorname{Spin}(16)$, see [9]. Its bracket is given by the above formula. ∎